Forecasting an Income Statement Assume Abercrombie & Fitch reports the following income statements. Income Statement, For Fiscal Years Ended ($ thousands) Feb. 2, 2013 Jan. 28, 2012 Jan, 29, 2011 Net sales $ 4,510,805 $ 4,158,058 $ 3,468,777 Cost of goods sold 1,694,096 1,607,834 1,251,348 Gross profit 2,816,709 2,550,224 2,217,429 Stores and distribution expense 1,987,926 1,888,248 1,589,501 Marketing, general and administrative expense 473,883 437,120 400,804 Other operating expense (income), net (19,333) 3,472 (10,056) Operating income 374,233 221,384 237,180 Interest expense, net 7,288 3,577 3,362 Income from continuing operations before taxes 366,945 217,807 233,818 Tax expense from continuing operations 129,934 74,669 78,109 Net income from continuing operations 237,011 143,138 155,709 Income from discontinued operations, net of tax -- 796 -- Net income $ 237,011 $ 143,934 $ 155,709 Forecast Abercrombie & Fitch’s fiscal 2014 income statement assuming the following income statement relations ($ in thousands). All percentages, other than sales growth and provision for income taxes, are based on percent of net sales. Net sales growth 1.10 Cost of goods sold/Net sales 37% Stores and distribution expense/Net sales 44% Marketing, general and administrative expense/Net sales 11% Other operating income, net $(19,333) Interest expense (income), net $7,288 Tax expense from continuing operations (% of pretax income) 35% Net loss from discontinued operations (net of taxes) 0 Round your answers to the nearest whole number. Do not use negative signs with your answers. Forecasted Income Statement, For Fiscal Years Ended ($ thousands) 2013 2014 Est. Net sales $ 4,510,805 $Answer 0 Cost of goods sold 1,694,096 Answer 0 Gross profit 2,816,709 Answer 0 Stores and distribution expense 1,987,926 Answer 0 Marketing, general and administrative expense 473,883 Answer 0 Other operating expense (income), net (19,333) Answer 0 Operating income 374,233 Answer 0 Interest expense, net 7,288 Answer 0 Income from continuing operations before taxes 366,945 Answer 0 Tax expense from continuing operations 129,934 Answer 0 Net income from continuing operations 237,011 Answer 0 Income from discontinued operations, net of tax -- Answer 0 Net income $ 237,011 $ Answer 0
Answer #1

Answer:

Forecasted Income Statement, For Fiscal Years Ended 2013 & 2014 ($ thousands)
2013 2014 Est. Calculation of 2014Est.
Net sales $4,510,805 $4,961,886 =4,510,805*1.1
Cost of goods sold           16,94,096 $1,835,898 =4,961,886*37%
Gross profit           28,16,709 $3,125,988
Stores and distribution expense           19,87,926 $2,183,230    =4,961,886*44%
Marketing, general and administrative expense             4,73,883 $545,807 =4,961,886*11%
Other operating expense (income), net               -19,333 ($19,333)
Operating income             3,74,233 $416,284
Interest expense, net                  7,288 $7,288
Income from continuing operations before taxes             3,66,945 $408,996
Tax expense from continuing operations             1,29,934 $143,149 =408,996*35%
Net income from continuing operations             2,37,011 $265,847 =408,996-143,149
Income from discontinued operations, net of tax                        -   $0
Net income $237,011 $265,847